# Revisiting the predictability of dynamical systems: a new local data-driven approach

### v1

@@ Appendix A Lorenz system and its stochastic version
where $dW^{j}_{t}$, $j=1,2,3$ indicate increments of three independent Wiener processes, and $\eta$ modulates the intensity of the stochastic perturbations. In Supplementary Figure 7, we show the predictability index for the Lorenz system with stochastic perturbations, using the same numerical simulation settings as in Figure 2, with the value of $\eta=1$. We note that the Supplementary Figure 7 is overall similar to Figure 2 for the five different forecasting horizons considered. However, we observe small-scale fluctuations of $\alpha_{\eta}$, which corresponds to the characteristics of the stochasticity within the system. In addition, the value of $\alpha_{\eta}$ drops more rapidly on average as expected, since the stochastic nature will make the system less predictable in general.
Figure 7: Distribution of $\alpha_{\eta}$ at different prediction horizons for the stochastic Lorenz-63 system. Same as Figure 2, but for the stochastic version of the Lorenz-63 system.
@@ Appendix B The scale-dependent nature of predictability
In this paper, all the results displayed in the main text use $q$ = 0.99, which means the closest one percent of states to the reference state are selected as recurrences. Figure 8 and 9 show the results for $q=0.98$ and $q=0.995$, respectively, following exactly the same settings as Figure 2. We note that using different $q$ values yields similar overall results, with the most prominent features being noticeable in each case. However, we also observe some differences: for larger q values ($q=0.995$) , small-scale stripes on the attractor are visible, whereas for relatively smaller q values, no such patterns exist and the variation of $\alpha_{\eta}$ in phase space appears smoother. This confirms the close relationship between predictability and scale, indicating that the choice of $q$ should be made according to specific application needs.
@@
In Figure 10, we further tested a wider range of $q$ values for their influence on the value of $\alpha_{\eta}$. Four states at different locations on the Lorenz-63 attractor are shown in Figure 10 (a), representing the wing, near the fixed point, before lobe transition (before the intersection of the two wings), and after lobe transition (after the intersection of the two wings). The $\alpha_{\eta}$ computed from $q$ ranging from 0.98 (2000 $recurrences$ for dataset with length of $10^{5}$) to 0.9999 (10 $recurrences$) are demonstrated in Figure 10 (b-e). We can see that although the overall differences given by different $q$ values are not significant, there are still noteworthy distinctions that confirm the scale-dependent nature of predictability.
@@
Figure 8: Comparative analysis of the impact of different quantile choices ($q$) on $\alpha_{t}$. Same as Fig. 2, but for $q=0.98$.
@@
Figure 9: Comparative analysis of the impact of different quantile choices ($q$) on $\alpha_{t}$. Same as Fig. 2, but for $q=0.995$.
@@
Figure 10: Predictability analysis for different $q$ values. (a) Illustration of the locations of the states considered in the Lorenz-63 phase space. (b-e) Heatmaps of $\alpha_{\eta}$ for four states located at the wing, near the fixed point, before lobe transition (before the intersection of the two wings), and after lobe transition (after the intersection of the two wings), respectively.
Appendix C Results for other dynamical systems
We also applied our predictability index to other idealized dynamical systems with varying complexity, including the Rössler system, and the spring-slider system for laboratory earthquakes.
C.1 Rössler system
$\frac{dx}{dt}$ $=$ $-y-z,$ (17)
$\frac{dy}{dt}$ $=$ $x+ay,$ (18)
$\frac{dz}{dt}$ $=$ $b+z(x-c).$ (19)
C.2 Spring-slider system for laboratory earthquakes
We follow the equations provided in [16] to generate the laboratory earthquakes dataset. We generated $2*10^{5}$ time steps with $dt=0.01$ for the analysis.
Figure 11: Distribution of $\alpha_{\eta}$ at different prediction horizons for the Rössler system. Same as Figure 2, but for the Rössler system and its corresponding Lyapunov time.
Figure 12: Distribution of $\alpha_{\eta}$ at different prediction horizons for the deterministic spring-slider system. Same as Figure 2, but for the deterministic spring-slider system.
Figure 13: Distribution of $\alpha_{\eta}$ at different prediction horizons for the stochastic spring-slider system. Same as Figure 2, but for the stochastic spring-slider system.
@@ Appendix D Comparison with Nonlinear Local Lyapunov Exponent
The Nonlinear Local Lyapunov Exponent (NLLE) was first proposed by Ding and Li in 2007 [36]. Unlike previous approaches that rely on linearized dynamics, NLLE uses the full dynamics to characterize the error growth rate without linearizing the governing equations. When the governing equations of the dynamical system is not explicitly known, they provided a method for estimating NLLE from observational data based on analogous, as shown in [82]. This index is sometimes also called the ’Local Growth Rate’, with an open-source implementation provided in [83]. In this section, we computed NLLE using the same Lorenz-63 dataset as used in the main analysis. Figure 14 demonstrates the distribution of NLLE ($\lambda_{\eta}$) for different $\eta$.
Figure 14: Distribution of NLLE at different prediction horizons for the stochastic Lorenz-63 system. Same as Figure 2, but for the Nonlinear Local Lyapunov Exponent.
@@ Appendix E Extension of the Predictability Index
In Figure 15 and 16, we show the results of ${\alpha_{\eta}^{\ast}}$ using $\phi(x)=x^{-1}$ and $\phi(x)=x$, respectively. For these two different $\phi(x)$, the results are generally similar to those obtained when $\phi(x)=1$, although they show noticeable differences in the details. When $\phi(x)=x^{-1}$, ${\alpha_{\eta}^{\ast}}$ decreases slowly as the forecasting horizon increases, whereas the opposite holds for $\phi(x)=x^{1}$. This can be explained as $\phi(x)=x^{-1}$ assigns higher weights to the close neighbors, while $\phi(x)=x^{1}$ assigns higher weights to the states that are farther away. This means we can selectively define $\phi$ based on the needs of different tasks. For instance, in some cases, we may be particularly interested in fast diverging trajectories, making $\phi(x)=x^{1}$ a more suitable choice.
@@
Figure 15: Distribution of ${\alpha_{\eta}^{\ast}}$ at different prediction horizons for the Lorenz-63 system, with $\phi(x)=x^{-1}$. Same as Figure 2, but for the extension of the predictability index.
@@
Figure 16: Distribution of ${\alpha_{\eta}^{\ast}}$ at different prediction horizons for the Lorenz-63 system, with $\phi(x)=x$. Same as Figure 2, but for the extension of the predictability index.
Definition of North Atlantic Wintertime Weather Regimes
@@
The definition of weather regimes is adapted from [74], focusing on the wintertime (DJFM) 500 hPa (Z500) in the Euro-Atlantic sector (80°W to 50°E, 22.5°N to 70°N). The weighted Z500 daily anomaly data is decomposed using EOF analysis, retaining the 14 leading modes (accounting for approximately 90% of the variance) to speed up the computation. The climatology is computed as the calendar day mean field of Z500 from 1979 to 2022, based on a centered 15-day window, while the weight is applied to the data as a cosine function of the latitude. The k-means clustering algorithm is then applied to the principal components, resulting in four weather regimes, whose centroids are demonstrated in Fig. 17. Next, we compute the normalized projection of each daily Z500 anomaly field onto the cluster centroid, which has been termed the Weather Regimes Index (IWR) in previous studies [84]. Here, we aim to analyze the predictability of established weather regimes; therefore, we require that the Weather Regimes Index (IWR) remains above 1.0 for at least 5 days to classify a state as belonging to a weather regime. States that are not allocated to any weather regime are classified as "no regime".
Real-time Predictability Analysis
@@
An example of a real-time application is presented using the Lorenz-63 system, following the same numerical simulation scheme as other experiments in this study. We assume that we have $10^{5}$ time steps as historical observations, and their corresponding $\alpha_{\eta}$ have already been computed, as shown in Fig. 18 (b,d,f). Meanwhile, we also assume there are 2000 states not present in the historical observations, which are used to simulate real-time unseen states in an operational forecast system. Figure 18 (a,c,e) shows the predictability proxy for the unseen states. By comparing the results with the reference $\alpha_{\eta}$, we note that the proposed proxy can effectively capture the features of predictability, regardless of the region in phase space or the forecasting horizon. Specifically, the low predictability near the lobe transition region is in Fig.18 (a,c), and the high predictability region near the fixed point in Fig.18 (e) are both well reflected. Although this application is a rather simple case, we anticipate its application in other operational forecast systems.
Figure 17: Composite Z500 anomaly maps for the four weather regimes in the North Atlantic-Europe region.

### v2

@@ Appendix A Lorenz system and its stochastic version
where $dW^{j}_{t}$, $j=1,2,3$ indicate increments of three independent Wiener processes, and $\eta$ modulates the intensity of the stochastic perturbations. In Fig. S1, we show the predictability index for the Lorenz system with stochastic perturbations, using the same numerical simulation settings as in Fig. 2, with the value of $\eta=1$. We note that the Fig. S1 is similar to Figure 2 for the five different forecasting horizons considered. However, we observe small-scale fluctuations of $\alpha_{\eta}$, which corresponds to the characteristics of the stochasticity within the system. In addition, the value of $\alpha_{\eta}$ drops more rapidly on average as expected, since the stochastic nature will make the system less predictable in general.
Figure S1: Distribution of $\alpha_{\eta}$ at different prediction horizons for the stochastic Lorenz-63 system. Same as Figure 2, but for the stochastic version of the Lorenz-63 system. The quantile $q$ applied in this analysis is 0.99, and the Theiler window size $w$ is set to 50 time steps.
@@ Appendix B The scale-dependent nature of predictability
In this paper, all the results displayed in the main text use $q$ = 0.99, which means the closest one percent of states to the reference state are selected as recurrences. Figure S2 and S3 show the results for $q=0.98$ and $q=0.995$, respectively, following exactly the same settings as Figure 2. We note that using different $q$ values yields similar overall results, with the most prominent features being noticeable in each case. However, we also observe some differences: for larger q values ($q=0.995$) , small-scale stripes on the attractor are visible, whereas for relatively smaller q values, no such patterns exist and the variation of $\alpha_{\eta}$ in phase space appears smoother. This confirms the close relationship between predictability and scale, indicating that the choice of $q$ should be made according to specific application needs.
@@
In Figure S4, we further tested a wider range of $q$ values for their influence on the value of $\alpha_{\eta}$. Four states at different locations on the Lorenz-63 attractor are shown in Fig. S4 (a), representing the wing, near the fixed point, before lobe transition (before the intersection of the two wings), and after lobe transition (after the intersection of the two wings). The $\alpha_{\eta}$ computed from $q$ ranging from 0.98 (2000 recurrencesrecurrences for dataset with length of $10^{5}$) to 0.9999 (10 recurrencesrecurrences) are demonstrated in Fig. S4 (b-e). We can see that although the overall differences given by different $q$ values are not significant, there are still noteworthy distinctions that confirm the scale-dependent nature of predictability.
@@
Figure S2: Comparative analysis of the impact of different quantile choices ($q$) on $\alpha_{t}$. Same as Fig. 2, but for $q=0.98$.
@@
Figure S3: Comparative analysis of the impact of different quantile choices ($q$) on $\alpha_{t}$. Same as Fig. 2, but for $q=0.995$.
@@
Figure S4: Predictability analysis for different $q$ values. (a) Illustration of the locations of the states considered in the Lorenz-63 phase space. (b-e) Heatmaps of $\alpha_{\eta}$ for four states located at the wing, near the fixed point, before lobe transition (before the intersection of the two wings), and after lobe transition (after the intersection of the two wings), respectively.
Appendix C Information theory-based extension of $\alpha_{\eta}$
In this section, we propose a link between the Shannon entropy, and our predictability index $\alpha_{\eta}$. Local predictability of a dynamical system can be interpreted as the divergence rate of neighboring trajectories. As shown in step 1 of Fig. S5, we can adopt the same definitions of $R_{t_{\zeta}}$ and $R_{t_{\zeta}}^{\eta}$ as in $\alpha_{\eta}$, and use second-order neighbors to characterize the divergence of states within $R_{t_{\zeta}}^{\eta}$ (step 2, in Fig. S5). Here, second-order neighbors are defined as the recurrences of all states within $R_{t_{\zeta}}$ and $R_{t_{\zeta}}^{\eta}$.
Let us set $R_{t_{\zeta}}$, and recall that it contains $N_{n}$ elements (states). Now, let us assume that for each of these elements, $N_{m}$ second-order neighbors are selected, where $N_{m}$ is a predefined parameter (set to $N_{n}$ in this section), resulting in a total of $N_{n}\times N_{m}$ potentially repeated states. These states can then be viewed as a probability distribution over all the states on the attractor by counting their occurrences and normalizing by $N_{n}\times N_{m}$, resulting in $p_{R_{t_{\zeta}}}$, as shown in Fig. S5. If the states in $R_{t_{\zeta}}$ are closely located in the phase space, the distribution will be more concentrated; conversely, if they are more spread, the distribution will be less concentrated. The same steps and analysis also apply to $R_{t_{\zeta}}^{\eta}$. If we now take the Shannon entropy perspective to characterize the predictability of a state $\zeta$}, we can compute the rate of change of the entropy over a time interval $\eta$: $\Delta H_{\eta}(\zeta)=H(p_{R_{t_{\zeta}}})-H(p_{R_{t_{\zeta}}^{\eta}})$ (step 3, in Fig. S5). This quantity measures the expected amount of information created. The more information is created, the more unexpected the configuration, and thus the less predictable the original state. We note that, in the illustrative example presented in Fig. S5, $\Delta H_{\eta}(\zeta)>0$, suggesting that $R_{t_{\zeta}}^{\eta}$ is more spread in the phase space than $R_{t_{\zeta}}$.
We consider this an effective link between $\alpha_{\eta}$ and information theory, noting that it would require further in-depth analysis, which lies beyond the scope of this paper. Yet, we believe that the proposed methodology may pave the way for new applications of information theory in high-dimensional dynamical systems.
Figure S5: Schematic illustration of the computation of entropy change $\Delta H_{\eta}(\zeta)$, demonstrated in the phase space of the Lorenz-63 system. The panels in the second row (step 1) provide a zoomed-in view of the phase space region where $\alpha_{\eta}(\zeta)$ is defined. Recurrences ($R_{t_{\zeta}}$) are represented by solid blue dots, while forward recurrences ($R_{t_{\zeta}}^{\eta}$) are depicted by empty blue dots, as in Fig. 1. The blue circle with radius $e^{-s(q,\zeta)}$ indicates the hypersphere used to define the neighborhood of the reference state. In step 2, second-order neighbors of each state within $R_{t_{\zeta}}$ and $R_{t_{\zeta}}^{\eta}$ are taken to construct probability distributions $p_{R_{t_{\zeta}}}$ and $p_{R_{t_{\zeta}}^{\eta}}$, respectively. The panels in the third row are colored according to these distributions. In step 3, the entropy of these two distributions is computed and compared.
@@ Appendix D Comparison with Nonlinear Local Lyapunov Exponent
The Nonlinear Local Lyapunov Exponent (NLLE) was first proposed by Ding and Li in 2007 [36]. Unlike previous approaches that rely on linearized dynamics, NLLE uses the full dynamics to characterize the error growth rate without linearizing the governing equations. When the governing equations of the dynamical system is not explicitly known, they provided a method for estimating NLLE from observational data based on analogous, as shown in [85]. This index is sometimes also called the ’Local Growth Rate’, with an open-source implementation provided in [86]. In this section, we computed NLLE using the same Lorenz-63 dataset as used in the main analysis. Figure S6 demonstrates the distribution of NLLE ($\lambda_{\eta}$) for different $\eta$.
Figure S6: Distribution of NLLE at different prediction horizons for the stochastic Lorenz-63 system. Same as Figure 2, but for the Nonlinear Local Lyapunov Exponent.
@@ Appendix E Extension of the Predictability Index
In Figure S7 and S8, we show the results of $\alpha_{\eta}^{\ast}$} using $\phi(x)=x^{-1}$ and $\phi(x)=x$, respectively. For these two different $\phi(x)$, the results are generally similar to those obtained when $\phi(x)=1$, although they show noticeable differences in the details. When $\phi(x)=x^{-1}$, $\alpha_{\eta}^{\ast}$} decreases slowly as the forecasting horizon increases, whereas the opposite holds for $\phi(x)=x^{1}$. This can be explained as $\phi(x)=x^{-1}$ assigns higher weights to the close neighbors, while $\phi(x)=x^{1}$ assigns higher weights to the states that are farther away. This means we can selectively define $\phi$ based on the needs of different tasks. For instance, in some cases, we may be particularly interested in fast diverging trajectories, making $\phi(x)=x^{1}$ a more suitable choice.
@@
Figure S7: Distribution of $\alpha_{\eta}^{\ast}$} at different prediction horizons for the Lorenz-63 system, with $\phi(x)=x^{-1}$. Same as Figure 2, but for the extension of the predictability index.
@@
Figure S8: Distribution of $\alpha_{\eta}^{\ast}$} at different prediction horizons for the Lorenz-63 system, with $\phi(x)=x$. Same as Figure 2, but for the extension of the predictability index.
Appendix F Definition of North Atlantic Wintertime Weather Regimes
@@
The definition of weather regimes is adapted from [63], focusing on the wintertime (DJFM) 500 hPa (Z500) in the Euro-Atlantic sector (80°W to 50°E, 22.5°N to 70°N). The weighted Z500 daily anomaly data is decomposed using EOF analysis, retaining the 14 leading modes (accounting for approximately 90% of the variance) to speed up the computation. The climatology is computed as the calendar day mean field of Z500 from 1979 to 2022, based on a centered 15-day window, while the weight is applied to the data as a cosine function of the latitude. The k-means clustering algorithm is then applied to the principal components, resulting in four weather regimes, whose centroids are demonstrated in Fig. S9. Next, we compute the normalized projection of each daily Z500 anomaly field onto the cluster centroid, which has been termed the Weather Regimes Index (IWR) in previous studies [87]. Here, we aim to analyze the predictability of established weather regimes; therefore, we require that the Weather Regimes Index (IWR) remains above 1.0 for at least 5 days to classify a state as belonging to a weather regime. States that are not allocated to any weather regime are classified as "no regime".
Figure S9: Composite Z500 anomaly maps for the four weather regimes in the North Atlantic-Europe region.
Appendix G Datesets for other examples
G.1 Rössler system
$\frac{dx}{dt}$ $=$ $-y-z,$ (21)
$\frac{dy}{dt}$ $=$ $x+ay,$ (22)
$\frac{dz}{dt}$ $=$ $b+z(x-c).$ (23)
G.2 Spring-slider system for laboratory earthquakes
In [16], one spring-slider model is developed to simulate the dynamics of slow earthquakes. This model is described by the following set of ordinary differential equations (ODEs):
$\dot{x}$ $=$ $\frac{e^{x}\left[\left(\beta_{1}-1\right)x(1+\lambda u)+y-u\right]+\kappa\left(\frac{v_{0}}{v_{*}}-e^{x}\right)-\dot{u}\frac{1+\lambda y}{1+\lambda u}}{1+\lambda u+ve^{x}}$ (24)
$\dot{y}$ $=$ $\kappa\left(\frac{v_{0}}{v_{*}}-e^{x}\right)-ve^{x}\dot{x}$ (25)
$\dot{z}$ $=$ $-\rho e^{x}\left(\beta_{2}x+z\right)$ (26)
$\dot{u}$ $=$ $-\alpha-\gamma u+\dot{z}.$ (27)
where the state of system can be fully characterized by the state vector $\left[x,y,z,u\right]$. To further simulate the irregularity of labquakes, two stochastic terms modeled by a Wiener process are added to the equations for $y$ and $u$, making the system a set of Stochastic Differential Equations (SDE):
$\mathrm{d}y$ $=$ $\left[\kappa\left(\frac{v_{0}}{v_{*}}-e^{x}\right)-ve^{x}\dot{x}\right]\mathrm{d}T+\varepsilon_{y}\mathrm{~d}W_{T}$ (28)
$\mathrm{~d}u$ $=$ $[-\alpha-\gamma u+\dot{z}]\mathrm{d}T+\varepsilon_{u}\mathrm{~d}W_{T}$ (29)
The values of all the parameters are presented in Table 1, while their physical explanations are available in [16]. We generated $2*10^{5}$ time steps with $dt=0.01$ for the analysis. In Figure 7, the attractor for the spring-slider system is visualized only only using variables x, y, and z, given the limitation to three dimensions.
Table 1: Simulation Parameters
Parameter Value
$\beta_{1}$ $1.2$
$\beta_{2}$ $0.8265$
$\lambda$ $0.0156$
$\kappa$ $0.2555$
$v_{0}$ $10^{-6}$
$v_{*}$ $10^{-6}$
$v$ $0.004$
$\rho$ $0.1$
$\alpha$ $0.0112$
$\gamma$ $0.03$
$\varepsilon_{y}$ $0.0115$
$\varepsilon_{u}$ $0.0111$
G.3 Double Pendulum, EEG and ECG
The data of these three systems are from [68], with the double pendulum data generated through numerical simulation, while the other two are obtained from observational biological data. The EEG and ECG data were adapted from the PhysioNet database [89, 90, 91].
For the double pendulum, the simulations were conducted using a variational integrator based on the Euler-Lagrange equations:
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{\partial L}{\partial\dot{\mathbf{q}}}-\frac{\partial L}{\partial\mathbf{q}}=0,$ (30)
with the Lagrangian L=TVL=T-V is the kinetic (T) minus potential (V) energy. For double pendulum, $\mathbf{q}=\left[\theta_{1}\quad\theta_{2}\right]^{T}$, and the Lagrangian (L) becomes:
$L=\frac{1}{2}(m_{1}+m_{2})l_{1}^{2}\dot{\theta}_{1}^{2}+\frac{1}{2}m_{2}l_{2}^{2}\dot{\theta}_{2}^{2}+m_{2}l_{1}l_{2}\dot{\theta}_{1}\dot{\theta}_{2}\cos(\theta_{1}-\theta_{2})-(m_{1}+m_{2})l_{1}g(1-\cos\theta_{1})-m_{2}l_{2}g(1-\cos\theta_{2}).$ (31)
where $l_{1}=l_{2}=m_{1}=m_{2}=1$, $g=10$. The action integral was approximated using the trapezoidal rule to compute the equations of motion:
$\delta\int_{a}^{b}L(\mathbf{q},\dot{\mathbf{q}},t)\,\mathrm{d}t=0.$ (32)
$x(t)=\sin(\theta_{1})$ is used as the observational time series for double pendulum.
For all these three examples, a similar approach to [68] is applied to the time series data to obtain eigen-time-delay coordinates. Specifically, a time series x(t) is stacked into a Hankel matrix H, which is then decomposed using singular value decomposition (SVD):
$\mathbf{H}=\left[\begin{array}[]{cccc}x\left(t_{1}\right)&x\left(t_{2}\right)&\cdots&x\left(t_{p}\right)\\ x\left(t_{2}\right)&x\left(t_{3}\right)&\cdots&x\left(t_{p+1}\right)\\ \vdots&\vdots&\ddots&\vdots\\ x\left(t_{q}\right)&x\left(t_{q+1}\right)&\cdots&x\left(t_{m}\right)\end{array}\right]=\mathbf{U}\Sigma\mathbf{V}^{*}.$ (33)
This yields a $\mathbf{V}^{*}$ matrix, which represents a hierarchy of eigen time series that reconstruct a delay-embedded attractor. The predictability index $\alpha_{\eta}$ is computed based on the $\mathbf{V}^{*}$ matrix and visualized on the attractor constructed from the three leading time series in $\mathbf{V}^{*}$ (see Fig. 7). A summary of the parameters used for the construction and decomposition of the Hankel matrix is provided in Table 2. We note that the parameters used here differ from those in [68] to retain higher energy for predictability analysis.
Table 2: Summary of Parameters
System Type Measured Variable Time Step $\Delta t$ # Rows in H $q$ Rank $r$ Energy in $r$ Modes (%)
Double Pendulum ODE $x(t)$ 0.001 100 5 99.9
ECG Data Voltage 0.004s 25 5 93.3
EEG Data Voltage 0.01s 200 8 61.2
Appendix H Real-time Predictability Analysis
@@
An example of a real-time application is presented using the Lorenz-63 system, following the same numerical simulation scheme as other experiments in this study. We assume that we have $10^{5}$ time steps as historical observations, and their corresponding $\alpha_{\eta}$ have already been computed, as shown in Fig. S10 (b,d,f). Meanwhile, we also assume there are 2000 states not present in the historical observations, which are used to simulate real-time unseen states in an operational forecast system. Figure S10 (a,c,e) shows the predictability proxy for the unseen states. By comparing the results with the reference $\alpha_{\eta}$, we note that the proposed proxy can effectively capture the features of predictability, regardless of the region in phase space or the forecasting horizon. Specifically, the low predictability near the lobe transition region is in Fig. S10 (a,c), and the high predictability region near the fixed point in Fig. S10 (e) are both well reflected. Although this application is a rather simple case, we anticipate its application in other operational forecast systems.